Calculate the critical radius (in centimeter) of a
Answer #1

a. \small \Delta T = 0.2T_{m}

crtical radius,

\small r^{*} = \frac{2\sigma _{sl}T_{m}}{\Delta H_{f}\Delta T}

substituting all the given values,

\small r^{*} = \frac{2\times 177\times 10^{-7}\times1083 }{1826\times 216.6}

\small r^{*} = 9.693 \times 10^{-8} cm

b. lattice parameter, a = 0.361 nm = 3.61 x 10-8 cm

for FCC, \small r = \frac{a\sqrt{2}}{4}

\small r = \frac{3.61\times 10^{-8}\times \sqrt{2}}{4}

\small r = 1.276 \times 10^{-8} cm

number of unit cell = r*/r = 7.596

number of atoms per unit cell of FCC = 4

number of atoms = 30.386